Please help! It looked easy to solve but I've been at it for fifty minutes and my solutions are not anywhere near (0,1). Can someone please help me? Thank You:)
Solve the system.
y=3^x
y=5^x
Thanks again!

Answer by dolly(163)

Hi,
given y = 3^x and y = 5^x
Here we equate both y
==> 3^x = 5^x
This equation is satisfied only when x = 0 and for no other value will this hold good as the bases are different.
==> 3^0 = 5^0
==> 1 = 1
Thus when x = 0, y = 3^0 = 5^0 = 1
So the only solution is (0, 1)
good luck!!!
dolly